OpenStudy (anonymous):

5x + y = 13 3x = 15 – 3y another system of equations problem

OpenStudy (ajprincess):

Divide the second eqn by 3 nd tell me vat u get?

OpenStudy (anonymous):

you mean multiply?

OpenStudy (anonymous):

5x + y = 13 3x = 15 – 3y Bring all x terms to one side for both equations, 5x + y = 13 3x - 13y = 15 We wish to make one of the coefficients same in both equations, so we can cancel it. 5x*3 + 3*y = 13*3 3x*5 + 13y*5 = 15*5 15x + 3y = 39 15x + 65y = 75 Cancel 15x from both equations, -62y = -36 62y = 36 y = 36/62 = 18/31